sophiaespinosa:

Loren solved the equation 10 = StartFraction 19 Over 9 EndFraction (149) + b for b as part of her work to find the equation of a trend line that passes through the points (1, 130) and (10, 149). What error did Loren make? A.) She should have solved 10 = StartFraction 9 Over 19 EndFraction (149) + b for b. B.)She should have solved 1 = StartFraction 19 Over 9 EndFraction (130) + b for b. C.)She should have solved 149 = StartFraction 19 Over 9 EndFraction (10) + b for b. D.)She should have solved 130 = StartFraction 9 Over 19 EndFraction (1) + b for b.

sophiaespinosa:

isk it would not let me do mathmatics

payton:

wym?

sophiaespinosa:

could you tell me what you are confused about

payton:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @sophiaespinosa isk it would not let me do mathmatics \(\color{#0cbb34}{\text{End of Quote}}\) the way you said this ^^

sophiaespinosa:

sorry i did the wrong key when writing idk and it would not let me pick the subject mathmatics

payton:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @sophiaespinosa sorry i did the wrong key when writing idk and it would not let me pick the subject mathmatics \(\color{#0cbb34}{\text{End of Quote}}\) thats weird...anyways...do you understand how to solve this question?

sophiaespinosa:

no

bruhman7624:

I cant tell you how hard I am dying from the comment before he said no XD

Mercury:

the generic equation for a line is y = mx + b using the equation they give you: 10 = (19/9)(149) + b it's clear that she plugged in (10, 149) into the equation y = (19/9)x + b and tried to solve for b. for the point (10,149), x = 10 and y = 149. but notice how this person reversed the x and y values (they plugged in y = 10 and x = 149 instead) with that being said, which answer choice shows them plugging in the numbers correctly?

sophiaespinosa:

A?

Mercury:

you're on the right track, but keep in mind answer A doesn't solve the problem that she switched the x and y values try swapping the x and y values in her work, and see what answer choice you get

sophiaespinosa:

D i'm sorry if i'm wrong i'm just panicking because its timed and I only have a little left

Mercury:

the x-value is 10, the y-value is 149 therefore, swapping 10 and 149 in the equation 10 = StartFraction 19 Over 9 EndFraction (149) + b gives you choice C

sophiaespinosa:

oh ok thank you i will look in on how you got the answer in the writing